# Discrete Messages Improve Communication Efficiency among Isolated Intelligent Agents

### v2

@@ Abstract
keywords:
discrete messages, multiple discrete tokens, inter-agent communication
1 Introduction
Intelligent agent communication, situated at the crossroads of AI and linguistics, aims to explore how a common language develops among agents. The Lewis Game[20] exemplifies collaborative tasks in this field, where a speaker and a listener collaborate to accomplish a task. The speaker describes a specific object to the listener, who must correctly identify it from a set of alternatives (Figure 1(left)). Using AI, significant research[31][16][35] has investigated language origins and evolution. Chaabouni et al[4] conducted research on emergent communication[15], focusing on various aspects such as the assessment of compositionality and generalization in emerging language agents and the development of efficient color naming systems[5]. Lazaridou et al[19] proposed a method that combines multi-agent communication with data-driven natural language learning, aiming to enable machine agents to effectively communicate with humans using natural language. Lowe et al[23] investigated the interaction between supervised learning and self-play in protocol learning for emergent communication. Deep reinforcement learning has also been applied in the field of multi-agent communication[30][7]. However, as the number of agents increases, redundant and indistinguishable information often leads to inefficient communication.Yet, with more agents comes the challenge of redundancy and inefficiency. Recent work[26][28][6][3] has delved into optimizing communication, surpassing simple capacity and extending dataset complexity, significantly contributing to our understanding of language’s evolution.
Vector Quantization (VQ) adheres to Shannon’s rate-distortion theory[9], suggesting that vector encoding may outperform scalar encoding by handling dependencies in source symbols adequately. Advances in reparameterization, specifically for VAEs handling discrete variables[29][24], have improved the model’s effectiveness. VQVAE[33] circumvents non-differentiability issues by applying the identity function for efficient gradient transmission. This vector-based discrete representation is shown to be more robust and generalizable than its continuous counterparts for complex learning models[33][22]. Furthermore, applying discretization in multi-agent reinforcement learning addresses communication challenges within modular reasoning architectures, enhancing efficient interactions across modules[11].
@@ 1 Introduction
Building upon the prior research, this study presents advancements in the vector quantization technique of the VQ model. We have made the following hypotheses in this paper: When intelligent agents have diverse personal experiences, using discrete messages is more advantageous than continuous messages, and communicating using multiple discrete tokens is more advantageous than using a single token. Through our designed experiments in multi-agent machine learning, we have empirically demonstrated that communication using sentences composed of multiple discrete tokens offers superior communication efficiency among agents with diverse personal experiences. However, in the experiments, when we use the VAE model to simulate continuous language communication between agents instead of the AE model, its effectiveness is superior to that of communication using multi-token discrete language.
Figure 1: Left: Lewis Game. Right: In structured architectures, communication is typically vectorized. The communication vector is initially divided into multiple discretization tokens. Each token goes through separate discretization, where it is quantized to the nearest neighbor within a shared collection of latent codebook vectors. Subsequently, the discretization tokens are concatenated back together to form a vector with the same shape as the original one.
@@
The structure of our paper is as follows: Section 1 provides the background of emergent communication and outlines the objectives of our research. In Section 2, we introduce the related work in the field. Section 3 describes the experimental settings and methods employed in our study. In Section 4, we present the experimental results and draw meaningful conclusions. Finally, in Section 5, we conclude our research with discussions and prospects for future work.
2 Related Works
In recent years, numerous approaches have emerged to facilitate effective communication among specialized components within machine learning models. These approaches include attention mechanisms that selectively transmit information between specialized components in machine learning models[11][2][10] and the Transform method[34][18]. Additionally, collective memory and shared parameter techniques have been utilized for communication in multi-agent settings[27].
The RIAL model is widely used for discrete communication in intelligent agent coordination, enabling communication through discrete symbols, mirroring human social interaction[7]. Guo et al.[12] explored the importance of computer simulations in evolutionary linguistics using intelligent agent models to simulate agents developing a compositional language for numerical concepts through communication. Studies into LSTM[14] language models by Lakretz et al.[17] have shed light on how hidden states represent numbers and syntactic structures, guiding further inquiry into language patterns. Further work by Miao et al.[25], Garcia et al.[8], and Havrylov et al.[13] expanded our understanding of multi-agent communication dynamics and the emergence of language in neural network-based agents.
In this vein, our research employs the Vector-Quantized Variational Autoencoder (VQVAE) model[33] and utilizes cross-training and cross-validation to probe communication patterns amongst agents. Our experiments demonstrate that in settings where agents have varied language systems, discrete language proves more effective than continuous forms. We also examine how differing token quantities in codebooks influence discrete communication effectiveness, a topic which is further elaborated in Section 4.
3 Theoretical Basis and Experimental Method
3.1 Discrete and Continuous Communication
In the series of autoencoder models[1], we designate the encoder component of the model as the speaker, while the decoder component is referred to as the listener.
@@ 3.1 Discrete and Continuous Communication
For the discrete communication model, we use VQVAE. To train a pair of agents, let’s assume the input is $x$. The information is passed through the speaker as $z_{e}(x)=e(x,\theta)$, representing the encoded representation of $x$. Then, the information undergoes discrete quantization using a codebook, resulting in $Z=DISCRETIZE(z_{e}(x),\varphi)$. Finally, the speaker reconstructs the original information from the received codebook indices $x^{{}^{\prime}}=d(Z,\phi)$, where the model parameters $\theta,\varphi,\phi$ are continuously updated by minimizing the reconstruction loss and codebook loss. The complete loss function for this process is as shown in Equation 1:
@@
Where the last two terms represent the codebook loss in the VQVAE model, in the subsequent algorithm, we use $L_{quantify}$ to represent these two items. The continuous data output by the encoder is quantized by the codebook layer before being transmitted to the decoder, which is the process of discrete communication between agents as we define it. In the experiments involving the AE model, the overall loss can be expressed as:$z=e(x,\theta)$,$x^{{}^{\prime}}=d(z,\phi)$. As shown in Equation 2, the overall loss is equivalent to the reconstruction loss.
@@
During this process, the continuous data output by the encoder is directly input into the decoder, which is the process of agents using continuous language for communication.
Throughout the entire experiment, the experimental data based on the Autoencoder (AE) serves as a baseline, which aims to verify that under the same experimental settings, the use of continuous language communication is less effective than discrete semantic communication between unfamiliar agents.
3.2 Multi-token Discretization
Li et al[21] proposed a human-like discrete information generation method that enables discrete message communication to have the effect of continuous message communication. Based on the foundation of discretization, we propose a multi-token discretization approach. In the original autoencoder (AE) framework, input data is encoded by the encoder into a continuous vector. The VQVAE model builds upon the AE and introduces a latent space codebook between the encoder and decoder. The continuous variables from the encoder are quantized into multiple vectors of the same size as the codebook. In our research, multi-token discretization is applied before the data enters the codebook layer. It involves dividing the output of the encoder into multiple segments of equal size but containing different data. Let’s assume our latent codebook size is $e\in R^{L\times M}$. Initially, the output $z_{e}(x)$ is divided into $N$ segments $s_{1},s_{2},s_{3},...s_{N}$ with $z_{e}(x)=CONCAT(s_{1},s_{2},s_{3},...s_{N})$, where each segment $s_{i}\in\mathbb{R}^{\frac{M}{N}}$ with $\frac{M}{N}\in N^{+}$. Next, each of these segments is discretized sequentially: $e_{o_{i}}=Discretize(s_{i})$, where $o_{i}=argmin||s_{i}-s_{j}||(j\in{1,....N})$. After the discretization process, the N segments of data are then integrated back together in the order of their original splitting, as shown in Equation 3:
$Z=CONCAT(Discretize(s_{1}),Discretize(s_{2}),...Discretize(s_{N}))$ (3)
Throughout the entire process, the discretized multi-token data always shares the same codebook. The schematic diagram of the multi-token discretization is illustrated in Figure 1(Right). After modifying this part of the structure, the corresponding model loss function needs to be adjusted as well. Since we divide the data into $N$ segments, the total loss function for model training is defined as shown in Equation 4.
@@ 3.2 Multi-token Discretization
Figure 2: Ten pairs of agents’ understanding of the same language. Left: Feature distribution of latent codebooks for different agents. Right: Similarity of different latent codebooks. The understanding of this language is different for each pair of agents.
3.3 Learning and Validation of Communication for Agents
@@
Attempts have been made to explore cross-training (Guo et al[12], Tieleman et al[32]) in the context of multi-agent learning. In this approach, during the simultaneous training of multiple agents, after each iteration, a random combination is selected, pairing one agent’s speaker with another agent’s listener for the next round of iterative learning. The reason behind this approach is that when multiple agents learn the same language, their understanding of the language may not be entirely identical. Figure 2 demonstrate the feature distributions of the latent codebook spaces for 10 pairs of agents trained simultaneously on the same MNIST dataset. Each color represents a communication protocol between a pair of agents, that is, the feature distribution in the codebook. It can be observed that there are differences in semantic understanding among the agents. Hence, cross-training becomes necessary because it allows different agents to have the most similar understanding of the same language. Algorithm 2 in the Appendix implements the aforementioned process.
Algorithm 1 Individual Training with Cross-Validation
1:Using the processed dataset: $Dataset^{{}^{\prime}}=(dataset_{1}^{{}^{\prime}},dataset_{2}^{{}^{\prime}},...dataset_{m}^{{}^{\prime}})$;
[…]
3:for each iteration $i$ do
4:     for each pair of agents $j$ do
5:         Sample input data $x_{j}$ from $dataset_{j}^{{}^{\prime}}$
6:         $L_{j}\Leftarrow d_{j}(c_{j}(e_{j}(x_{j})))$
7:         optimize $e_{j}$,$c_{j}$ and $d_{j}$ with respect to $L_{j}$
8:     end for
[…]
10:for each pair of agents $j$ do
11:     Sample validate data $x$
12:     $loss_{j}\Leftarrow d_{k}(c_{j}(e_{j}(x))),(k=1,2...m,k\neq j)$
13:     output $loss_{j}$
14:end for
@@
In terms of experimental validation, in addition to using trained agents for verification, we conducted another form of validation by manipulating the dataset. Assuming there are $m$ types of data in the dataset, we merged the training set and validation set into a single dataset. The merged dataset was then divided into $m$ classes based on their labels $Dataset=(dataset_{1},dataset_{2},...dataset_{m})$. After the division, a portion of images was uniformly sampled from each class to form the validation set, denoted as $Validataset=(sample_{1}^{1},sample_{2}^{1},...sample_{m}^{1})$. To meet the experimental overlap requirements, from the remaining training set $P_{train}$, a certain number of images were extracted from each class according to the desired experimental overlap rate across the classes $Overlapset=(sample_{1}^{2},sample_{2}^{2},...sample_{m}^{2})$. There are an equal number of images in each class, and $P_{train}$ is the number of images left after the first extraction. For $p_{j}\in\left\{0.05,0.1,0.2,...0.9\right\}$, the calculation of the number of images sampled in the second extraction is given by Equation 5.
@@
These extracted images were then merged with the respective training sets, ensuring that each class in the training set contained images from the remaining $m-1$ classes. The processed training set is $Dataset^{{}^{\prime}}=(dataset_{1}^{{}^{\prime}},dataset_{2}^{{}^{\prime}},...dataset_{m}^{{}^{\prime}})$. This process resulted in a training set where each category served as a separate training set for a pair of agents to learn from. The experiments conducted on the split dataset follow Algorithm 1, which forms the core of our research paper. Specifically, discrete messages are more effective than continuous ones when agents have diverse personal experiences. Similar to Algorithm 1, the experimental methodology of our core content is illustrated in Figure 3.
The aforementioned are the two experimental procedures we used to explore the communication patterns of multiple agents. The main model involved in the procedures is the VQVAE model. However, when we incorporate the AE (Autoencoder) model in our experiments, we simply remove the ”c” module from the procedure, and the data outputted by the encoder is directly decoded by the decoder.
Figure 3: Training and validation of paired agents. Each pair of agents has its own dataset during training and learning. Upon completion of learning, one agent from the pair interacts with the other agent from another pair. This is our core methodology, where in this validation scenario, the advantage of discrete language in communication between agents is determined based on the reconstruction losses of information.
4 Experiments
The models used in this paper were implemented on the PyTorch 1.12.1 framework, using PyCharm Community Edition 2023.1 on the Windows platform. The model training was conducted on a single GeForce RTX 3060 GPU with 8GB of GPU memory, using the CUDA 12.1 experimental environment. The operating system used was Windows. In our work, we employed three datasets: MNIST, CIFAR10, CelebA and our own medical dataset. The image resolution for all three datasets was separately set to $28\times 28,32\times 32,64\times 64$ and $64\times 64$. The batch size for the first two datasets during training is set to 256, while the batch size for the latter two datasets is set to 64, and we utilized the Adam optimizer with a learning rate of 0.001. The commitment cost for the model’s discrete layer was set to 0.25, with a decay rate of 0.99. The specific codebook size for the discrete layer varied depending on the dataset. In the experiments, we compared the reconstruction errors between the original images and the reconstructed images. An example of the two types of images can be seen in Figure 4.
@@ 4 Experiments
4.1 Multi-token Discretization for Improved Agent Communication
@@
According to the method shown in Figure 3 and Algorithm 1, we repeated the experiments with different overlap ratios using the multi-token discretization model with the best performance and the AE model. For the MNIST and CelebA datasets, the original VQVAE model had latent space size $e_{m}\in\mathbb{R}^{512\times 64}$ and $e_{m}\in\mathbb{R}^{512\times 128}$, while for the CIFAR10 dataset, it was $e_{m}\in\mathbb{R}^{1024\times 256}$. We conducted the above overlap experiments with 32 tokens, and the experimental results are shown in Figure 5. Under three different datasets, the average loss incurred by using multiple discrete tokens for communication is 32.1%, 10.6%, and 3.7% lower than that incurred by using continuous semantics for communication, respectively. Our experiments indicate that when one agent interacts with another unfamiliar agent, the discrete semantic learning method using multiple tokens has certain advantages over continuous semantic learning.
Figure 5: Communication test loss on three types of models. The bar graphs above represent the losses for $m$ pairs of agents interacting with continuous variables, discrete variables, and multi-token discrete variables under different dataset overlap probabilities. The communication between agents that learn and interact with multi-token discrete variables is the most effective.
@@
Figure 6 explain why we chose to conduct our experiments with a 32-token VQVAE model as mentioned above and also illustrate the advantages of our proposed multi-token discrete mechanism compared to a single-token approach. It shows the results of training $m$ pairs of agents simultaneously according to Algorithm 1, where each boxplot in the figure represents the stable loss from communications between the $m$ pairs of agents. The general pattern is that as the number of discrete tokens increases, the communication loss decreases, which from the model’s perspective, leads to improved performance, and in terms of agent communication, this makes their communication more efficient.
@@
Figure 6: The communication loss of multiple agents under multi-token discretization. The increase in the number of tokens can reduce the loss from communications. Left: MNIST, codebook size $e_{m}\in\mathbb{R}^{512\times 64}$. Mid: CIFAR10, codebook size $e_{m}\in\mathbb{R}^{1024\times 256}$. Right: CelebA, codebook size $e_{m}\in\mathbb{R}^{512\times 64}$.
@@
In all of the above experiments, the number of agent pairs $m$ for the three datasets respectively are ($m=10(MNIST,CIFAR10),8(CelebA)$). Our experiments have demonstrated two theories. First, discrete messages with multiple tokens are more effective than continuous ones when agents have diverse personal experiences. Second, communications using multiple discrete tokens are more advantageous than those using a single token.
4.2 Theoretical Validation and Practical Application

### v3

@@ Abstract
Index Terms:
Discrete tokens, inter-agent communication, agents
I Introduction
Intelligent agent communication, positioned at the intersection of AI and linguistics, explores the development of a shared language among agents. The Lewis Game[18] serves as a key example of collaborative tasks in this area, where a speaker and listener work together to identify a specific object from a set of alternatives (Figure 1(left)). This field has seen extensive research into the origins and evolution of language through AI, with studies investigating various aspects of emergent communication. [4] focused on the compositionality and generalization capabilities of language agents and the creation of efficient color naming systems[5]. [17] proposed blending multi-agent communication with data-driven natural language learning to facilitate machine-human interaction. [21] explored the interplay between supervised learning and self-play in developing communication protocols for emergent communication, while deep reinforcement learning has been applied to tackle the challenges of multi-agent communication[28, 7], especially as the number of agents increases, leading to potential redundancy and inefficiency in communication. Recent research[24, 26, 6, 3] has focused on optimizing communication strategies, expanding our understanding of language evolution by addressing both capacity and dataset complexity challenges. Vector Quantization (VQ), in line with Shannon’s rate-distortion theory[9], suggests that vector encoding can outperform scalar encoding by effectively handling dependencies in source symbols. Recent advancements in reparameterization, particularly for VAEs managing discrete variables[27, 22], have enhanced model effectiveness. The VQ-VAE[30] model overcomes non-differentiability issues by employing the identity function for efficient gradient transmission, proving to be more robust and generalizable than continuous counterparts for complex learning models[30, 20]. Moreover, the application of discretization in multi-agent reinforcement learning tackles communication challenges within modular reasoning architectures, facilitating efficient interactions across modules[11].
@@ I Introduction
This study builds on previous research by advancing the vector quantization technique in VQ models. Our hypotheses suggest that when intelligent agents have diverse personal experiences, communicating via discrete messages, especially through sentences composed of multiple discrete tokens, is more effective than using continuous messages. Our experiments in multi-agent machine learning have empirically demonstrated that communication through multi-token discrete sentences significantly enhances communication efficiency among agents with diverse experiences. However, when employing the VAE model to simulate continuous language communication between agents, rather than the AE model, its effectiveness surpasses that of multi-token discrete language communication.
@@
The structure of our paper is organized as follows: Section I provides a background on agents’ communication and the objectives of our research. Section II introduces related work in the field. Section III details the experimental settings and methods employed in our study. In Section IV, we present our experimental results and draw meaningful conclusions. Finally, Section V concludes our research with discussions and prospects for future work.
Figure 1: Left: Lewis Game. Right: Multi-token Discrete Mechanism. The communication vector is initially divided into multiple discretization tokens. Each token goes through separate discretization, where it is quantized to the nearest neighbor within a shared collection of latent codebook vectors. Subsequently, the discretization tokens are concatenated back together to form a vector with the same shape as the original one.
II Related Work
In recent years, several methods have been devised to improve communication within machine learning models, notably through attention mechanisms[11, 2, 10] and the Transform method[31, 16]. Furthermore, collective memory and shared parameters have enhanced multi-agent communication[25]. The Reinforced Inter-Agent Learning (RIAL) model stands as a prominent framework for discrete communication among intelligent agents, supporting the use of discrete symbols to enable interactions reminiscent of human social behaviors[7]. The work of [12] delves into the crucial role of computer simulations in evolutionary linguistics, illustrating how intelligent agent models can foster the development of compositional languages for numerical concepts through communication. Research into LSTM[14] language models by [15] has illuminated the manner in which hidden states encapsulate numerical values and syntactic structures, spurring further exploration into linguistic patterns. Subsequent studies[23, 8, 13] have expanded our comprehension of the dynamics of multi-agent communication and the genesis of language among neural network-based agents.
Building on this foundation, our research leverages the Vector-Quantized Variational Autoencoder (VQ-VAE) model[30] and adopts cross-training and cross-validation techniques to scrutinize communication patterns between agents. Our findings reveal that in environments where agents employ diverse language systems, discrete forms of language are more efficacious than continuous ones. We also delve into how the variation in token numbers within codebooks affects the efficiency of discrete communication, a subject that receives in-depth treatment in Section IV.
III Theoretical Basis and Experimental Method
In the series of autoencoder models[1], we designate the encoder component of the model as the speaker, while the decoder component is referred to as the listener. The communication between the speaker and the listener within the same model is considered internal communication, while the communication between the speaker or listener and other model components is referred to as agents’ communication. We define $e\left(\cdot\right)$ as the encoder function, $d\left(\cdot\right)$ as the decoder function, and $h\left(\cdot\right)$ as the quantization layer function of VQ-VAE. The codebook space of the quantization layer is denoted as $C=\left[c_{1},c_{2},...c_{n}\right]$.
III-A Discrete and Continuous Communication
@@
For the discrete communication model, we use VQ-VAE. To train a pair of agents, let’s assume the input is $x$. The information is passed through the speaker as $z_{e}(x)=e(x,\theta)$, representing the encoded representation of $x$. Then, the information undergoes discrete quantization using a codebook, resulting in $Z=h(z_{e}(x),\varphi)$. Finally, the speaker reconstructs the original information from the received codebook indices $x^{{}^{\prime}}=d(Z,\phi)$, where the model parameters $\theta,\varphi,\phi$ are continuously updated by minimizing the reconstruction loss and codebook loss. The complete loss function for this process is as shown in Equation 1:
@@
Where the last two terms represent the quantization loss in the VQ-VAE model, in the subsequent algorithm, we use $L_{quantify}$ to represent these two items. In the experiments involving the AE model, the overall loss can be expressed as:$z=e(x,\theta)$,$x^{{}^{\prime}}=d(z,\phi)$. As shown in Equation 2, the overall loss is equivalent to the reconstruction loss.
@@
During this process, the continuous data output by the encoder is directly input into the decoder, which is the process of agents using continuous language for communication. And the involvement of the codebook quantization layer mentioned above refers to the process of discrete communication.
Throughout the entire experiment, the experimental data based on the Autoencoder (AE) serves as a baseline, which aims to verify that under the same experimental settings, the use of continuous communication is less effective than discrete communication between unfamiliar agents.
III-B Multi-token Discretization
[…]
@@ III-B Multi-token Discretization
Figure 2: Ten agents’ understanding of the same language. Left: Feature distribution of latent codebooks for different agents. Right: Similarity of different latent codebooks. The understanding of this language is different for each agent.
III-C Learning and Validation of Communication for Agents
@@
Attempts have been made to explore cross-training[12, 29] in the context of multi-agent learning. In this approach, during the simultaneous training of multiple agents, after each iteration, a random combination is selected, pairing one agent’s speaker with another agent’s listener for the next round of iterative learning. The reason behind this approach is that when multiple agents learn the same language, their understanding of the language may not be entirely identical. Figure 2 demonstrate the feature distributions of the latent codebook spaces for $10$ agents trained simultaneously on the same MNIST dataset. Each color represents a internal communication protocol, that is, the feature distribution in the codebook. It can be observed that there are differences in semantic understanding among agents. Hence, cross-training becomes necessary because it allows different agents to have the most similar understanding of the same language. Algorithm 2 in the Appendix A implements the aforementioned process.
Algorithm 1 Individual Training and communication
1:Using the processed dataset: $train_{-}set=(train_{-}set_{1},train_{-}set_{2},...train_{-}set_{m})$
2:Train $j$ agents simultaneously
3:Initialize encoders $E=\{e_{0}\left(\cdot\right),...,e_{m}\left(\cdot\right)\}$.
4:Initialize quantization layers $H=\{h_{0}\left(\cdot\right),...,h_{m}\left(\cdot\right)\}$
5:Initialize decoders $D=\{d_{0}\left(\cdot\right),...,d_{m}\left(\cdot\right)\}$
6:for each iteration $i$ do
7:     for each agent $j$ do
8:         Sample input data $x_{j}$ from $train_{-}set_{j}$
9:         $L_{j}\Leftarrow MSE(x_{j},d_{j}(h_{j}(e_{j}(x_{j})))$
10:         optimize $e_{j}\left(\cdot\right)$,$h_{j}\left(\cdot\right)$ and $d_{j}\left(\cdot\right)$ with respect to $L_{j}$
11:     end for
12:end for
13:Validation:
14:for each agent $j$ do
15:     Sample validate data $x$ from $val_{-}set$
16:     $loss_{j}\Leftarrow MSE(x,d_{k}(h_{j}(e_{j}(x)))),(k=1,2...m,k\neq j)$
17:     output $loss_{j}$
18:end for
19:$Communication\frac{}{}loss=\sum loss_{j}$
@@
In terms of experimental validation, in addition to using trained agents for verification, we conducted another form of validation by manipulating the dataset. Assuming there are $m$ types of data in the dataset, we merged the training set and validation set into a single dataset. The merged dataset was then divided into $m$ classes based on their labels $Dataset=(dataset_{1},dataset_{2},...dataset_{m})$. After the division, a portion of images was uniformly sampled from each class to form the validation set, denoted as $val_{-}set=(sample_{1}^{1},sample_{2}^{1},...sample_{m}^{1})$. To meet the experimental overlap requirements, from the remaining training set $P_{train}$, a certain number of images were extracted from each class according to the desired experimental overlap rate across the classes $Overlapset=(sample_{1}^{2},sample_{2}^{2},...sample_{m}^{2})$. There are an equal number of images in each class, and $P_{train}$ is the number of images left after the first extraction. For $p_{j}\in\left\{0.05,0.1,0.2,...0.9\right\}$, the calculation of the number of images sampled in the second extraction is given by Equation 4.
@@
$Overlapset$ were then merged with the respective training sets $dataset_{i}^{{}^{\prime}}$, where $dataset_{i}^{{}^{\prime}}=dataset_{i}-sample_{i}^{1}-sample_{i}^{2}$, ensuring that each class in the training set contained images from the remaining $m-1$ classes. The processed training set is $train_{-}set=(train_{-}set_{1},train_{-}set_{2},...train_{-}set_{m})$. This process resulted in a training set where each category served as a separate training set for single agent to learn from. The experiments conducted on the split dataset follow Algorithm 1, which forms the core of our research paper. Similar to Algorithm 1, the experimental methodology of our core content is illustrated in Figure 3.
The aforementioned are the two experimental procedures we used to explore the communication patterns of multiple agents. The main model involved in the procedures is the VQ-VAE model. However, when we incorporate the AE model in our experiments, we simply remove quantization layers $H$ from the procedure, and the data outputted by the encoder is directly decoded by the decoder.
Figure 3: Training and validation of agents. Each agent has its own dataset during training. Upon completion of learning, one agent interacts with the other agent . This is our core methodology, where in this validation scenario, the advantage of discrete language in communication between agents is determined based on the reconstruction losses of information.
IV Experiments
In our work, we employed four datasets: MNIST, CIFAR10, CelebA and Diabetic Retinopathy dataset. The image resolution for all four datasets was separately set to $28\times 28,32\times 32,64\times 64$ and $64\times 64$. The batch size for the first two datasets during training is set to $256$, while the batch size for the latter two datasets is set to $64$, and we utilized the Adam optimizer with a learning rate of $0.001$. The commitment cost for the model’s discrete layer was set to $0.25$, with a decay rate of $0.99$. The specific codebook size for the discrete layer varied depending on the dataset. In the experiments, we evaluate the effectiveness of agents’ communication by measuring the error between the original images and the reconstructed images. An example of the two types of images can be seen in Figure 4.
@@ IV Experiments
IV-A Multi-token Discretization for Improved Agents’ Communication
@@
According to the method shown in Figure 3 and Algorithm 1, we repeated the experiments with different overlap ratios using the multi-head discretized VQ-VAE model with the best performance and the AE model. For the MNIST and CelebA datasets, the original VQ-VAE model had latent space size $C_{m}\in\mathbb{R}^{512\times 64}$ and $C_{m}\in\mathbb{R}^{512\times 128}$, while for the CIFAR10 dataset, it was $C_{m}\in\mathbb{R}^{1024\times 256}$. We conducted the above overlap experiments with $32$ tokens, and the experimental results are shown in Figure 5. Under three different datasets, the average loss incurred by using multiple discrete tokens for communication is 32.1%, 10.6%, and 3.7% lower than that incurred by using continuous semantics for communication, respectively. Our experiments indicate that when one agent interacts with another unfamiliar agent, the discrete semantic learning method using multiple tokens has certain advantages over continuous semantic learning.
Figure 5: Communication loss on three types of models. The x-axis represents different overlap ratios, and the y-axis represents the communication loss between agents.
@@
Figure 6 explain why we chose to conduct our experiments with a 32-token VQ-VAE model and also illustrate the advantages of our proposed multi-token discrete mechanism compared to a single-token approach. It shows the results of training $m$ agents simultaneously according to Algorithm 1, where each boxplot in the figure represents the stable loss from communications between the $m$ agents. The general pattern is that as the number of discrete tokens increases, the communication loss decreases.
@@
Figure 6: The communication loss of multiple agents with the same data overlap ratio under multi-token discretization. The increase in the number of tokens can reduce the loss from communications.
@@
In all of the above experiments, the number of agents $m$ for the three datasets respectively are ($m=10(MNIST,CIFAR10),8(CelebA)$). Our experiments have demonstrated two theories. First, discrete communication with multiple tokens are more effective than continuous ones when agents have diverse personal experiences. Second, communications using multiple discrete tokens are more advantageous than those using a single token.
IV-B Theoretical Validation and Practical Application